# ReGenesis: LLMs can grow into Reasoning generalists via Self-Improvement

### v1

@@ Abstract
Post-training Large Language Models (LLMs) with explicit reasoning trajectories can enhance their reasoning abilities. However, acquiring such high-quality trajectory data typically demands meticulous supervision from humans or superior models, which can be either expensive or license-constrained. In this paper, we explore how far an LLM can improve its reasoning by self-synthesizing reasoning paths as training data without any additional supervision. Existing self-synthesizing methods, such as STaR, suffer from poor generalization to out-of-domain (OOD) reasoning tasks. We hypothesize it is due to that their self-synthesized reasoning paths are too task-specific, lacking general task-agnostic reasoning guidance. To address this, we propose Reasoning Generalist via Self-Improvement (ReGenesis), a method to self-synthesize reasoning paths as post-training data by progressing from abstract to concrete. More specifically, ReGenesis self-synthesizes reasoning paths by converting general reasoning guidelines into task-specific ones, generating reasoning structures, and subsequently transforming these structures into reasoning paths, without the need for human-designed task-specific examples used in existing methods. We show that ReGenesis achieves superior performance on all in-domain and OOD settings tested compared to existing methods. For six OOD tasks specifically, while previous methods exhibited an average performance decrease of approximately $4.6\%$ after post training, ReGenesis delivers around $6.1\%$ performance improvement. We also conduct in-depth analysis of our framework and show ReGenesis is effective across various LLMs and design choices.
@@ 1 Introduction
To fill this gap, in this paper we explore how to self-synthesize reasoning paths as post training data to make LLMs generalize well across various OOD reasoning tasks. We hypothesize that existing methods fail to generalize to OOD tasks because their synthesized reasoning paths are overly tailored to specific training tasks and lack diversity. These methods predominantly utilize a chain-of-thought approach, which leads to uniform and narrow reasoning patterns, rather than employing a range of cognitive strategies that foster diverse thinking steps. To resolve these shortcomings, we introduce ReGenesis, a reasoning generalist framework designed to generate general and diverse reasoning paths guided by various reasoning principles. ReGenesis adopts a methodological approach that transitions from abstract and general reasoning guidelines to concrete and specific reasoning solutions, as shown in Figure 1. More specifically, we first create multiple general task-agnostic reasoning guidelines in the beginning of the reasoning path generation. Then, given each general reasoning guideline, we prompt the LLM to augment such general guideline into a task-specific reasoning guideline. Note that in this step we fill in task-specific details into general guidelines while still maintaining their task-agnostic reasoning structure, as shown in Figure 1. Then we feed each task-specific reasoning guideline into the given LLM to generate a reasoning structure without directly solving the tasks. Each reasoning structure is subsequently used to create one candidate reasoning path. Finally, all the candidate the reasoning paths are filtered using ground truth answers and, along with the instructions and answers, are used to fine-tune the same LLM, enhancing its reasoning capabilities.
@@ 3.2 Filtering Reasoning Paths
Adding Ground-truth as a Hint. For instruction $q_{i}$ that does not have any correct reasoning solutions after filtering, we include the ground-truth $a_{i}$ as a hint to generate adapted guidance and reasoning structures (system prompts can be found in Appendix A.3.3). Unlike Zelikman et al. (2022) which provides hints directly in the final reasoning path generation process, we only add hints during “Guidance Adaption” step and “Reasoning Structure Generation” step. By not having ground-truth answers in the final reasoning generation stage, we reduce the risk of the final reasoning solutions producing a correct answer but having an incorrect reasoning path. We show the effectiveness of this design choice in Section 5.3. Additionally, incorporating ground truth as a hint significantly improved the utilization of the training set, increasing the average usage from $96.2\%$ to $99.4\%$ across $5$ datasets, as shown in Table 1.
@@ 4.2 In-Domain Experiments
During our experiments, we also noted that training solely with ground-truth answers (FT w/ GT) and excluding reasoning paths does not consistently improve the performance of the base large language model (LLM), even on in-domain test sets. Specifically, in the GSM8K and ARC-C tasks, relying exclusively on ground-truth answers led to performance declines of 44.8% and 3.4%, respectively. This observation underscores the importance of incorporating reasoning paths during the fine-tuning stage to enhance the LLM’s reasoning capabilities.
@@
Test Datsets
Training Training ASDIV SVAMP AQUA BBH ANLI OpenbookQA
Datasets Methods (Math) (Math) (Math) (Logical) (NLI) (Commonsense)
w/o FT $77.2\%$ $75.4\%$ $41.3\%$ $60.8\%$ $38.4\%$ $75.6\%$
FT w/ GT $54.0\%$downarrow $40.0\%$downarrow $29.1\%$downarrow $53.6\%$downarrow $44.6\%$uparrow $72.6\%$downarrow
GSM8K LMSI w/ GT $77.3\%$uparrow $72.2\%$downarrow $31.1\%$downarrow $59.5\%$downarrow $43.4\%$uparrow $73.2\%$downarrow
(Math) STaR $79.6\%$uparrow $71.5\%$downarrow $46.9\%$uparrow $47.4\%$downarrow $45.0\%$uparrow $72.8\%$downarrow
Ours $\mathbf{81.2\%}$uparrow $\mathbf{83.9\%}$uparrow $\mathbf{48.8\%}$uparrow $\mathbf{69.3\%}$uparrow $\mathbf{49.5\%}$uparrow $\mathbf{81.4\%}$uparrow
w/o FT $77.2\%$ $75.4\%$ $41.3\%$ $60.8\%$ $38.4\%$ $75.6\%$
FT w/ GT $53.8\%$downarrow $54.9\%$downarrow $32.3\%$downarrow $42.4\%$downarrow $37.0\%$downarrow $64.8\%$downarrow
NumGLUE LMSI w/ GT $75.7\%$downarrow $78.2\%$uparrow $40.6\%$downarrow $59.5\%$downarrow $35.1\%$downarrow $72.6\%$downarrow
(Math) STaR $\mathbf{79.6\%}$uparrow $76.1\%$uparrow $37.0\%$downarrow $58.5\%$downarrow $41.9\%$uparrow $71.6\%$downarrow
Ours $76.9\%$downarrow $\mathbf{79.4\%}$uparrow $\mathbf{48.4\%}$uparrow $\mathbf{61.7\%}$uparrow $\mathbf{50.0\%}$uparrow $\mathbf{79.8\%}$uparrow
w/o FT $77.2\%$ $75.4\%$ $41.3\%$ $60.8\%$ $38.4\%$ $75.6\%$
FT w/ GT $62.0\%$downarrow $56.0\%$downarrow $22.0\%$downarrow $44.6\%$downarrow $47.1\%$uparrow $71.4\%$downarrow
ReClor LMSI w/ GT $77.9\%$uparrow $76.8\%$uparrow $45.6\%$uparrow $53.9\%$downarrow $35.1\%$downarrow $72.4\%$downarrow
(Logical) STaR $76.1\%$downarrow $74.6\%$downarrow $46.1\%$uparrow $60.6\%$downarrow $42.5\%$uparrow $77.2\%$uparrow
Ours $\mathbf{76.4\%}$downarrow $\mathbf{76.5\%}$uparrow $\mathbf{49.6\%}$uparrow $\mathbf{66.8\%}$uparrow $\mathbf{44.8\%}$uparrow $\mathbf{81.4\%}$uparrow
w/o FT $77.2\%$ $75.4\%$ $41.3\%$ $60.8\%$ $38.4\%$ $75.6\%$
FT w/ GT $57.7\%$downarrow $57.2\%$downarrow $18.9\%$downarrow $37.6\%$downarrow $36.3\%$downarrow $78.4\%$uparrow
ARC-c LMSI w/ GT $70.9\%$downarrow $72.0\%$downarrow $32.7\%$downarrow $60.8\%$ $32.5\%$downarrow $79.4\%$uparrow
(Logical) STaR $77.0\%$downarrow $76.2\%$uparrow $40.6\%$downarrow $60.7\%$downarrow $\mathbf{47.4\%}$uparrow $\mathbf{84.2\%}$uparrow
Ours $\mathbf{81.6\%}$uparrow $\mathbf{79.5\%}$uparrow $\mathbf{46.5\%}$uparrow $\mathbf{66.0\%}$uparrow $46.4\%$uparrow $82.8\%$uparrow
w/o FT $77.2\%$ $75.4\%$ $41.3\%$ $60.8\%$ $38.4\%$ $75.6\%$
FT w/ GT $69.4\%$downarrow $72.3\%$downarrow $43.7\%$uparrow $52.3\%$downarrow $44.9\%$uparrow $62.2\%$downarrow
StrategyQA LMSI w/ GT $56.3\%$downarrow $56.8\%$downarrow $40.6\%$downarrow $60.6\%$downarrow $39.8\%$uparrow $68.4\%$downarrow
(Commonsense) STaR $79.8\%$uparrow $76.2\%$uparrow $\mathbf{43.3\%}$uparrow $62.9\%$uparrow $37.3\%$downarrow $77.4\%$uparrow
Ours $\mathbf{81.3\%}$uparrow $\mathbf{81.1\%}$uparrow $42.9\%$uparrow $\mathbf{65.9\%}$uparrow $\mathbf{55.3\%}$uparrow $\mathbf{80.4\%}$uparrow
@@ 5.1 Choices of Design
Test Datsets
Training Ablation GSM8K ASDIV SVAMP AQUA BBH ANLI OpenbookQA
Datasets Set (Test) (Math) (Math) (Math) (Logical) (NLI) (Commonsense)
A+P $66.2\%$ $76.1\%$ $80.3\%$ $47.6\%$ $53.4\%$ $42.1\%$ $75.6\%$
GSM8K S+P $70.0\%$ $77.5\%$ $80.1\%$ $47.2\%$ $57.9\%$ $43.0\%$ $76.4\%$
(Train) A+S+P $76.6\%$ $81.2\%$ $83.9\%$ $48.8\%$ $69.3\%$ $49.5\%$ $81.4\%$
@@ 5.2 Filtering Methods
5.3 Different Positions to add hint
In this experiment, we validated the hypothesis presented in Section 3.2, demonstrating that removing ground-truth answers in the final reasoning generation stage reduces the likelihood of producing correct answers with incorrect reasoning paths. Unlike ReGenesis, which first generates reasoning paths without hints and then adds hints only to those instructions lacking correct solutions, here we applied hints to all instructions in the GSM8K (Cobbe et al., 2021) training set for generating reasoning paths. To evaluate the impact of adding hints on performance, we compared two approaches: adding hints during “Guidance Adaptation” and “Reasoning Structure Generation” (A+S) versus adding ground-truth answers at every step including the “Reasoning Path Generation (A+S+P). As shown in Table 6 , the A+S approach outperformed A+S+P, showing better results on in-domain tests and 5 out of 6 OOD tests. This experiment also confirmed that the reasoning paths generated with hints were of lower quality than those generated without them, supporting our approach of generating reasoning paths first and adding hints only for instructions without correct solutions in ReGenesis.
@@ 5.4 Effect of the language model
Test Datsets
Training Hint GSM8K ASDIV SVAMP AQUA BBH ANLI OpenbookQA
Datasets Position (Test) (Math) (Math) (Math) (Logical) (NLI) (Commonsense)
GSM8K A+S $\mathbf{75.7\%}$ $\mathbf{79.3\%}$ $\mathbf{79.7\%}$ $\mathbf{50.0\%}$ $67.3\%$ $\mathbf{41.9\%}$ $\mathbf{77.4\%}$
(Train) A+S+P $73.7\%$ $78.8\%$ $78.4\%$ $44.1\%$ $\mathbf{69.0\%}$ $41.3\%$ $77.0\%$
@@
Test Datsets
Training Training NumGLUE ASDIV SVAMP AQUA BBH ANLI OpenbookQA
Datasets Methods (Test) (Math) (Math) (Math) (Logical) (NLI) (Commonsense)
w/o FT $58.9\%$ $88.2\%$ $86.8\%$ $\mathbf{64.6\%}$ $36.5\%$ $53.4\%$ $82.2\%$
NumGLUE STaR $75.6\%$ $88.3\%$ $87.9\%$ $61.0\%$ $69.4\%$ $52.3\%$ $81.2\%$
(Train) ReGenesis $\mathbf{79.2\%}$ $\mathbf{89.2\%}$ $\mathbf{91.1\%}$ $63.4\%$ $\mathbf{70.1\%}$ $\mathbf{53.9\%}$ $\mathbf{83.8\%}$
@@
5.5 Diverse Preferences for Reasoning Guidelines
As detailed in Section 3.1, ReGenesis  converts $25$ general seed reasoning guidelines into $25$ different reasoning paths for each instruction within the dataset. We are curious to see if the effectiveness of these guidelines vary across different base models. Therefore, for two different models, the Mistral-7B-Instruct-v0.3 and Meta-Llama-3-8B-Instruct models, we apply ReGenesis to the 25 guidelines for each instruction in the NumGLUE (Mishra et al., 2022) dataset and compute the average success rate ($x_{i}$) for each guideline across the entire dataset. To facilitate comparison of guideline preferences between models, we perform z-score normalization ($z=\frac{x-\mu}{\sigma}$, where $\mu$ and $\sigma$ are mean and standard deviation of the success rate $x=\{x_{1},....,x_{25}\}$ ) on the success rates and select the $10$ guidelines with the largest discrepancies between the two models. As shown in Figure 2, different language models vary in their preference for guidelines when self-synthesizing reasoning paths. For instance, Mistral-7B-Instruct-v0.3 demonstrates a preference for “7. Use systems thinking: Consider the problem as part of a larger system and understanding the interconnectedness of various elements.”. In contrast, Meta-Llama-3-8B-Instruct favors “21. Let’s think step by step.”. It also suggests that existing methods, such as STaR, which generate reasoning paths as post-training data, might overlook these preferences in their designed CoT examples. In contrast, ReGenesisautomates the process, leading to superior OOD performance.
@@ Appendix A Appendix
A.1 Diversity of the Training Dataset
@@
A.2 Hyperparameters of the self-synthesized dataset creation
@@
A.3 Details of Reasoning Generalist
A.3.1 Seed Prompts
@@
A.3.2 Prompts used for generating reasoning paths without hints
@@
A.3.3 Prompts used for generating reasoning paths with hints
@@
A.4 Details of Training
A.4.1 Baseline: Fine-Tuning with Ground Truth (FT w/ GT)
@@
A.4.2 Baseline: Fine-Tuning with Self-Improvement approach (LMSI)
@@
A.4.3 Baseline: Fine-Tuning with Self-Taught Reasoner approach (STaR)
@@ A.4.4 ReGenesis
A.5 Training details of Section 5
A.5.1 Training details: DIVERSITY OF THE TRAINING DATASET
@@
A.5.2 Training details: EFFECT OF THE LANGUAGE MODEL

### v2

@@ Abstract
Post-training Large Language Models (LLMs) with explicit reasoning trajectories can enhance their reasoning abilities. However, acquiring such high-quality trajectory data typically demands meticulous supervision from humans or superior models, which can be either expensive or license-constrained. In this paper, we explore how far an LLM can improve its reasoning by self-synthesizing reasoning paths as training data without any additional supervision. Existing self-synthesizing methods, such as STaR, suffer from poor generalization to out-of-domain (OOD) reasoning tasks. We hypothesize it is due to that their self-synthesized reasoning paths are too task-specific, lacking general task-agnostic reasoning guidance. To address this, we propose Reasoning Generalist via Self-Improvement (ReGenesis  [^1] The code is available at: https://github.com/SalesforceAIResearch/ReGenesis) , a method to self-synthesize reasoning paths as post-training data by progressing from abstract to concrete. More specifically, ReGenesis self-synthesizes reasoning paths by converting general reasoning guidelines into task-specific ones, generating reasoning structures, and subsequently transforming these structures into reasoning paths, without the need for human-designed task-specific examples used in existing methods. We show that ReGenesis achieves superior performance on all in-domain and OOD settings tested compared to existing methods. For six OOD tasks specifically, while previous methods exhibited an average performance decrease of approximately $4.6\%$ after post training, ReGenesis delivers around $6.1\%$ performance improvement. We also conduct in-depth analysis of our framework and show ReGenesis is effective across various LLMs and design choices.
@@ 1 Introduction
To fill this gap, we explore how to self-synthesize reasoning paths as post training data to make LLMs generalize well across various OOD reasoning tasks. We hypothesize that existing methods fail to generalize to OOD tasks because their synthesized reasoning paths are overly tailored to specific training tasks and lack diversity. These methods predominantly utilize a chain-of-thought approach, which leads to uniform and narrow reasoning patterns, rather than employing a range of cognitive strategies that foster diverse thinking steps. To resolve these shortcomings, we introduce ReGenesis, a reasoning generalist framework designed to generate general and diverse reasoning paths guided by various reasoning principles. ReGenesis adopts a methodological approach that transitions from abstract and general reasoning guidelines to concrete and specific reasoning solutions, as shown in Figure 1. More specifically, we first create multiple general task-agnostic reasoning guidelines in the beginning of the reasoning path generation. Then, given each general reasoning guideline, we prompt the LLM to augment such general guideline into a task-specific reasoning guideline. Note that in this step we fill in task-specific details into general guidelines while still maintaining their task-agnostic reasoning structure, as shown in Figure 1. Then we feed each task-specific reasoning guideline into the given LLM to generate a reasoning structure without directly solving the tasks. Each reasoning structure is subsequently used to create one candidate reasoning path. Finally, all the candidate the reasoning paths are filtered using ground truth answers and, along with the instructions and answers, are used to fine-tune the same LLM, enhancing its reasoning capabilities.
@@ 3.2 Filtering Reasoning Paths
Adding Ground-truth as a Hint. For instruction $q_{i}$ that does not have any correct reasoning solutions after filtering, we include the ground-truth $a_{i}$ as a hint to generate adapted guidance and reasoning structures (system prompts can be found in Appendix A.4.3). Unlike Zelikman et al. (2022) which provides hints directly in the final reasoning path generation process, we only add hints during “Guidance Adaption” step and “Reasoning Structure Generation” step. By not having ground-truth answers in the final reasoning generation stage, we reduce the risk of the final reasoning solutions producing a correct answer but having an incorrect reasoning path. We show the effectiveness of this design choice in Appendix A.1. Additionally, incorporating ground truth as a hint significantly improved the utilization of the training set, increasing the average usage from $96.2\%$ to $99.4\%$ across $5$ datasets, as shown in Table 1.
@@ 4.2 In-Domain Experiments
During our experiments, we also noted that training solely with ground-truth answers (FT w/ GT) and excluding reasoning paths does not consistently improve the performance of the base large language model (LLM), even on in-domain test sets. Specifically, in the GSM8K and ARC-C tasks, relying exclusively on ground-truth answers led to performance declines of 44.8% and 3.4%, respectively. This observation underscores the importance of incorporating reasoning paths during the fine-tuning stage to enhance the LLM’s reasoning capabilities. We also conduct an analysis to identify the reasons behind the significant performance on GSM8K in the Appendix A.8.1. The results reveal that the significant improvement is driven by the diverse answering strategies, which are perfectly aligned with the specific needs of GSM8K.
@@
[…]
@@ 5.1 Choices of Design
[…]
@@ 5.3 Effect of the language model
[…]
@@
5.4 Diverse Preferences for Reasoning Guidelines
As detailed in Section 3.1, ReGenesis converts $25$ general seed reasoning guidelines into $25$ different reasoning paths for each instruction within the dataset. We are curious to see if the effectiveness of these guidelines vary across different base models. Therefore, for two different models, the Mistral-7B-Instruct-v0.3 and Meta-Llama-3-8B-Instruct models, we apply ReGenesis to the 25 guidelines for each instruction in the NumGLUE (Mishra et al., 2022) dataset and compute the average success rate ($x_{i}$) for each guideline across the entire dataset. To facilitate comparison of guideline preferences between models, we perform z-score normalization ($z=\frac{x-\mu}{\sigma}$, where $\mu$ and $\sigma$ are mean and standard deviation of the success rate $x=\{x_{1},....,x_{25}\}$ ) on the success rates and select the $10$ guidelines with the largest discrepancies between the two models. As shown in Figure 2, different language models vary in their preference for guidelines when self-synthesizing reasoning paths. For instance, Mistral-7B-Instruct-v0.3 demonstrates a preference for “7. Use systems thinking: Consider the problem as part of a larger system and understanding the interconnectedness of various elements.”. In contrast, Meta-Llama-3-8B-Instruct favors “21. Let’s think step by step.”. It also suggests that existing methods, such as STaR, which generate reasoning paths as post-training data, might overlook these preferences in their designed CoT examples. In contrast, ReGenesisautomates the process, leading to superior OOD performance. Besides LLMs, we also demonstrate in the Appendix A.8.3 how different training datasets exhibit varying preferences for different seed reasoning guidelines.
@@ 5.4 Diverse Preferences for Reasoning Guidelines
5.5 Preservation of Task-Agnostic Reasoning in Task-Specific Paths
In this experiment, we demonstrated that although ReGenesis uses the final detailed and task-specific reasoning paths as training data, the diverse task-agnostic reasoning strategies that ReGenesis adopts to generate such final task-specific paths, are effectively preserved in the final task-specific paths such that it is easy for LLMs, such as Mistral-7B-Instruct-v0.3, to detect them and learn them from the final task-specific paths. To evaluate this, we randomly select $500$ reasoning paths from ReGenesis and STaR’s generated data of NumGLUE training data set. We then prompt Mistral-7B-Instruct-v0.3 to label each of the total 1000 reasoning paths the task-agnostic reasoning strategy behind the task-specific path. Next, we task Mistral-7B-Instruct-v0.3 to cluster the task-agnostic reasoning strategies generated by itself and suggest cluster names for both ReGenesis and STaR. For STaR, we obtain a small number of clusters with limited diversity, such as “Direct Calculation and Simplification” and “Problem Decomposition and Step-by-Step Calculation”. In contrast, ReGenesis produces a significantly more diverse set of clusters, such as “Questioning Assumptions and Critical Thinking”, “Systematic Thinking”, and “Reflective Thinking”. More detailed results can be found in Appendix A.8.2. Results show that it is quite apparent for Mistral-7B-Instruct-v0.3 that the final task-specific paths generated by ReGenesis, have more diverse task-agnostic reasoning strategies behind them, compared to STaR. We have also run the same analysis at OOD inference on OpenBookQA, on models fine-tuned with ReGenesis and STaR separately on NumGLUE dataset. Similar results can be found in Appendix A.8.2. Our findings indicate that models trained with ReGenesis, actually learned to adopt more diverse task-agnostic reasoning strategies at inference in OOD settings.
@@ Appendix A Appendix
A.1 Different Positions to add hint
In this experiment, we validated the hypothesis presented in Section 3.2, demonstrating that removing ground-truth answers in the final reasoning generation stage reduces the likelihood of producing correct answers with incorrect reasoning paths. Unlike ReGenesis, which first generates reasoning paths without hints and then adds hints only to those instructions lacking correct solutions, here we applied hints to all instructions in the GSM8K (Cobbe et al., 2021) training set for generating reasoning paths.
[…]
To evaluate the impact of adding hints on performance, we compared two approaches: adding hints during “Guidance Adaptation” and “Reasoning Structure Generation” (A+S) versus adding ground-truth answers at every step including the “Reasoning Path Generation (A+S+P). As shown in Table 7 , the A+S approach outperformed A+S+P, showing better results on in-domain tests and 5 out of 6 OOD tests. This experiment also confirmed that the reasoning paths generated with hints were of lower quality than those generated without them, supporting our approach of generating reasoning paths first and adding hints only for instructions without correct solutions in ReGenesis.
A.2 Diversity of the Training Dataset
@@
A.3 Hyperparameters of the self-synthesized dataset creation
@@
A.4 Details of Reasoning Generalist
A.4.1 Seed Prompts
@@
A.4.2 Prompts used for generating reasoning paths without hints
@@
A.4.3 Prompts used for generating reasoning paths with hints
@@
A.5 Details of Training
A.5.1 Baseline: Fine-Tuning with Ground Truth (FT w/ GT)
@@
A.5.2 Baseline: Fine-Tuning with Self-Improvement approach (LMSI)
@@
A.5.3 Baseline: Fine-Tuning with Self-Taught Reasoner approach (STaR)
@@ A.5.4 ReGenesis
A.6 Training details of Section 5
A.6.1 Training details: DIVERSITY OF THE TRAINING DATASET
@@
A.6.2 Training details: EFFECT OF THE LANGUAGE MODEL